1. -3 = 7 - 4y
2. 2x + 4(3x - 6) = 4 - (6x + 2)
HELP, HELP

Answer by minerva(22)

wel heres ur answer
1. -3 = 7 - 4y
1. -3-7=-4y
-10=-4y
y=10/4
y=2.5
2. 2x + 4(3x - 6) = 4 - (6x + 2)
2. 2x+12x -24=4-6x-2
14x+6x=4-2+24
20x=26
x=26/20
x= 1 6/20